OpenStudy (anonymous):

{(7 3/8)(23 1/5)} / (7 1/4)

OpenStudy (saifoo.khan):

\[7\frac{3}{8}*23\frac{1}{5} \div 7\frac{1}{4}\]

OpenStudy (saifoo.khan):

this?

OpenStudy (anonymous):

Yes

OpenStudy (saifoo.khan):

\[23\frac{3}{5} \]

OpenStudy (anonymous):

22 4/5 was what it said the answer was now I have to do another problem for the question. {(6 7/8)(21 3/5)} / (7 3/4)

OpenStudy (anonymous):

Here is an example it shows but I have no idea how they get this answer. {(7 1/8)(23 3/5)}/(7 1/2) = 22 2/5

OpenStudy (anonymous):

When I get the answer wrong it just gives the same problem with different numbers and you are able to solve it until you get it right

OpenStudy (saifoo.khan):

are u doing Mathlab?

OpenStudy (anonymous):

I have no idea how they are getting these answers and yes

OpenStudy (saifoo.khan):

thts the problem with online schools.

OpenStudy (saifoo.khan):

Im using my calc for them. but i also knw how to do them manually. first convert them to improper frac, then solve them!

OpenStudy (saifoo.khan):

what's the next prob?

OpenStudy (anonymous):

{(6 7/8)(21 3/5)} / (7 3/4)

OpenStudy (anonymous):

55/8 * 108/5 / 31/4? is that right?

OpenStudy (anonymous):

5940/40 * 4/31?

OpenStudy (saifoo.khan):

right.

OpenStudy (anonymous):

23760/1240 = 19.161...... then I am lost how do I turn that into a fraction?

OpenStudy (saifoo.khan):

try cancelling the frac by 2.

OpenStudy (anonymous):

like this? 11880/620 = 1188/62 = 594/31 = that can't be right can it

OpenStudy (saifoo.khan):

u r doing it right.

OpenStudy (anonymous):

2376/124 = 594/31 = 19 5/31? then rounded to the nearest fifth would be

OpenStudy (saifoo.khan):

right.

OpenStudy (anonymous):

It has to be rounded to the nearest fifth

OpenStudy (saifoo.khan):

19.161290 what fifth? deicaml? sign. fig?

OpenStudy (anonymous):

fraction

OpenStudy (saifoo.khan):

how can it be a nearest to fifth fraction?

OpenStudy (anonymous):

like 19 1/5 I am not sure that's what it says to do

OpenStudy (saifoo.khan):

Nope. that must be in decimals.

OpenStudy (anonymous):

well it's a word problem about measuring head circumference for a hat and it says to round to the nearest 1/5th of an inch

OpenStudy (anonymous):

so there's what 32 cm in an inch right or 32 something

OpenStudy (saifoo.khan):

aaah, i did that before! its in Mathlab.. there are same 2 problems!

OpenStudy (saifoo.khan):

i wrote that in decialms that time.